# Shedding Light on Gravity: Black Hole Shadows and Lensing Signatures in Lorentz Gauge Theory

### v1

@@ V Shadow of the Black Hole
The photon sphere around a black hole is defined by unstable circular photon orbits that are significantly influenced by the gravitational field. Photons in these orbits either spiral into the black hole’s event horizon or escape to infinity, forming a luminous photon ring. This ring delineates the observable boundary of the black hole shadow for distant observers Synge (1966); Cunningham and Bardeen (1972); Bardeen et al. (1972); Luminet (1979a). Historically, different terminologies such as escape cone Synge (1966), cone of gravitational radiation capture Zeldovich and Novikov (1971), optical appearance of black holes, and black hole image Bardeen (1973); Chandrasekhar (1998); Luminet (1979b, 2018) have been utilized. However, the contemporary and widely adopted term black hole shadow, introduced by Falcke et al. Falcke et al. (2000), refers explicitly to the dark region surrounded by the photon ring Johannsen and Psaltis (2010c); Johnson et al. (2020).
In this section, we investigate the properties of the black hole shadow, focusing particularly on its dependence on the underlying model parameters, such as $A_{0}$, $m_{s}$, and $m_{p}$. Given the spherical symmetry of the spacetime considered, the analysis is simplified by restricting null geodesic trajectories to the equatorial plane ($\theta=\pi/2$). The metric in this case reduces to:
@@
$\boxed{r_{\mathrm{sh}}=\frac{3\sqrt{3}A_{0}^{3}m_{s}}{8\pi m_{p}^{2}}}.$ (68)
Considering a Schwarzschild black hole scenario (with $m_{s}=1$) and a known shadow radius $r_{\mathrm{sh}}=5.2$, we derive the parameter relationship:
$\frac{A_{0}^{3}}{m_{p}^{2}}=\frac{8\pi r_{\mathrm{sh}}}{3\sqrt{3}}\approx 25.164.$ (69)
This relation constrains the combination $A_{0}^{3}/m_{p}^{2}$, yet it allows flexibility in selecting either parameter individually. For example:
$A_{0}\approx 2.93\quad\text{if}\quad m_{p}=1,$ (70)
$m_{p}\approx 0.199\quad\text{if}\quad A_{0}=1.$ (71)
Figures 3 and 4 highlight the sensitivity of the photon sphere and shadow radii to the parameter $A_{0}$. Both radii exhibit monotonic growth with increasing $A_{0}$. The shadow radius, however, demonstrates a notably stronger dependence due to its cubic relationship with $A_{0}$, indicating how alterations in $A_{0}$ significantly affect the geometry around the black hole. This result underscores the potential observational impact of different model parameters in gravitational lensing phenomena.
Figure 3: Shadow radius and photon sphere radius as functions of the parameter $A_{0}$, assuming $m_{p}=m_{s}=1$.
Figure 4: Visualization of the black hole shadow with an accretion disk, highlighting the geometrical effect of parameter variations ($m_{p}=m_{s}=1$).
@@ VI Strong lensing by the black hole
$A_{0}\dot{t}_{0}^{2}=C_{0}\dot{\phi}_{0}^{2},$ (72)
where $X_{0}\equiv X(r_{0})$. Using the definition of the impact parameter in Eq. (62) at $r=r_{0}$, the function $R(r)$ in Eq. (63) can be rewritten as
$R(r)=\frac{A_{0}C}{AC_{0}}-1.$ (73)
We define the critical impact parameter as Bozza (2002b)
$b_{c}\equiv b(r_{\mathrm{ph}})=\lim_{r_{0}\rightarrow r_{\mathrm{ph}}}\sqrt{\frac{C_{0}}{A_{0}}}.$ (74)
In Ref. Bozza (2002b), the deflection angle of light rays passing at the radial distance $r_{0}$ from a static asymptotically flat black hole is given as
[…]
where $R(r)$ is given in Eq. (73). Introducing the variable $z\equiv 1-r_{0}/r$, the integral in Eq. (75) can be separated into the divergent part, $I_{D}(r_{0})$, and the regular part, $I_{R}(r_{0})$. The divergent part is expressed as Bozza (2002b); Tsukamoto (2018); Ali et al. (2025)
[…]
@@
$x_{1}(r_{0})=\frac{1-A_{0}}{C_{0}A_{0}^{\prime}}\left(C_{0}^{\prime}A_{0}-C_{0}A_{0}^{\prime}\right),$ (78a)
$x_{2}(r_{0})=\frac{(1-A_{0})^{2}}{2C_{0}^{2}A_{0}^{\prime 3}}\Big{[}2C_{0}C_{0}^{\prime}A_{0}^{\prime 2}+(C_{0}C_{0}^{\prime\prime}-2C_{0}^{\prime 2})A_{0}A_{0}^{\prime}-C_{0}C_{0}^{\prime}A_{0}A_{0}^{\prime\prime}\Big{]}.$ (78b)
@@
$x_{1}(r_{0})=\frac{C_{\mathrm{ph}}r_{\mathrm{ph}}\mathfrak{L^{\prime}_{\mathrm{ph}}}}{B_{\mathrm{ph}}}\left(r_{0}-r_{\mathrm{ph}}\right)+\mathcal{O}(r_{0}-r_{\mathrm{ph}})^{2},$ (80)
@@
$b(r_{0})=b_{c}+\frac{1}{4}\sqrt{\frac{C_{\mathrm{ph}}}{A_{\mathrm{ph}}}}\,\mathfrak{L}_{\mathrm{ph}}^{\prime}\left(r_{0}-r_{\mathrm{ph}}\right)^{2}+\mathcal{O}(r_{0}-r_{\mathrm{ph}})^{2}.$ (81)
@@
$\lim_{r_{0}\rightarrow r_{\mathrm{ph}}}x_{1}(r_{0})=\lim_{b\rightarrow b_{c}}\frac{2C_{\mathrm{ph}}r_{\mathrm{ph}}\sqrt{\mathfrak{L}_{\mathrm{ph}}^{\prime}}}{B_{\mathrm{ph}}}\left(\frac{b}{b_{c}}-1\right)^{1/2}.$ (82)
@@
$I_{D}(b)=-\frac{r_{\mathrm{ph}}}{\sqrt{x_{2}(r_{\mathrm{ph}})}}\ln\left(\frac{b}{b_{c}}-1\right)+\frac{r_{\mathrm{ph}}}{\sqrt{x_{2}(r_{\mathrm{ph}})}}\ln\Bigl{(}r_{\mathrm{ph}}^{2}\mathfrak{L}^{\prime}_{\mathrm{ph}}\Bigr{)}+\mathcal{O}(b-b_{c}),$ (83)
@@
$x_{2}(r_{\mathrm{ph}})=\frac{C_{\mathrm{ph}}(1-A_{\mathrm{ph}})^{2}\Bigl{[}C_{\mathrm{ph}}^{\prime\prime}A_{\mathrm{ph}}-C_{\mathrm{ph}}A^{\prime\prime}_{\mathrm{ph}}\Bigr{]}}{2A_{\mathrm{ph}}^{2}C_{\mathrm{ph}}^{\prime}}.$ (84)
@@
$\bar{a}=\sqrt{\frac{2B_{\mathrm{ph}}A_{\mathrm{ph}}}{C^{\prime\prime}_{\mathrm{ph}}A_{\mathrm{ph}}-C_{\mathrm{ph}}A^{\prime\prime}_{\mathrm{ph}}}},$ (87a)
$\bar{k}=\bar{a}\ln\Bigl{(}r_{\mathrm{ph}}^{2}\mathfrak{L}_{\mathrm{ph}}^{\prime}\Bigr{)}+I_{R}(r_{\mathrm{ph}})-\pi.$ (87b)
@@
$b_{c}=\sqrt{\frac{C_{\mathrm{ph}}}{A_{\mathrm{ph}}}}=\frac{3\sqrt{3}A_{0}^{3}m_{S}}{8\pi m_{P}^{2}},$ (88)
@@
Since this integral cannot be computed analytically for general $A_{0}$, we note that for the Schwarzschild black hole case ($A_{0}=\pm 1$), the integral evaluates to $I_{R}^{\mathrm{SBH}}\approx 0.9496$, which matches the result reported in Ref. Bozza (2002b). More generally, as $A_{0}^{2}\to 1$, the values of $I_{R}$ approach this Schwarzschild black hole value (see Fig. 5).
@@
Additionally, for $A_{0}^{2}=1$, we find $\bar{k}=-0.40023$, which corresponds to the Schwarzschild black hole lensing result reported in Ref. Bozza (2002b). Finally, by incorporating these relations, we present in Fig. 6 the $b$-profile of the strong deflection angle as a function of $A_{0}$.
@@ VI.1 The lens equation and the EHT constraints
Assuming that $D_{OL}\gg b_{c}$, the angular position of the $n$th relativistic image is given by Bozza (2002c)
@@
To apply these theoretical results to astrophysical observations, we consider the supermassive black holes M87* and Sgr A*. The black hole M87* has a mass of $(6.5\pm 0.7)\times 10^{9}\,M_{\odot}$ and is located at a distance of $D_{OL}=16.8\,\mathrm{Mpc}$ from Earth [The Event Horizon Telescope Collaboration] et al. (2019a, b). Meanwhile, Sgr A* has a mass of $4^{+1.1}_{-0.6}\times 10^{6}\,M_{\odot}$ and is situated at a distance of $7.97\,\mathrm{kpc}$ from Earth [The Event Horizon Telescope Collaboration] et al. (2022a); Akiyama et al. (2022). Substituting these parameters into Eq. (99), Fig. 7 illustrates the outermost Einstein rings ($n=1$) for M87* and Sgr A*, assuming they are black holes within the framework of the LGT. These rings are depicted in the celestial coordinate system of an observer on Earth, with coordinates $X$ and $Y$, and are computed for different values of $A_{0}$.
@@
A key quantity in strong lensing is the magnification of the $n$th relativistic image, defined as Virbhadra and Ellis (2000c); Bozza (2002c)
@@
Since the magnification is inversely proportional to $D_{OL}^{2}$, relativistic images are typically faint. The outermost image appears the brightest, while the brightness of higher-order images diminishes exponentially. However, in the limiting case where $\psi$ approaches zero, implying near-perfect alignment of the source and the lens, the images undergo significant magnification. It is noteworthy that while the outermost image, corresponding to $\theta_{1}$, remains distinct, higher-order images tend to cluster around $\theta_{\infty}\equiv\theta_{n}|_{n\to\infty}$ Bozza (2002c), which is identified as $\theta_{\infty}=\theta_{c}$, representing the innermost relativistic image.
@@
while the second is the relative magnification between the outermost relativistic image and the collective group of inner relativistic images, expressed as Bozza (2002c)
@@
In 2019, the Event Horizon Telescope (EHT) collaboration published the first horizon-scale image of the supermassive black hole M87*, providing strong observational evidence for black holes. Their analysis revealed that the compact emission region had an angular diameter of $\theta_{d}=(42\pm 3)\,\mu\text{as}$, along with a central flux depression exceeding a factor of $\gtrsim 10$, corresponding to the black hole’s shadow [The Event Horizon Telescope Collaboration] et al. (2019c, a, b). Subsequently, in 2022, the EHT collaboration captured an image of Sgr A*, the supermassive black hole at the center of the Milky Way. This image displayed a distinctive ring structure with an angular diameter of $\theta_{d}=(48.7\pm 7)\,\mu\text{as}$ and a deviation from the Schwarzschild shadow characterized by $\delta=-0.08^{+0.09}_{-0.09}$ (VLTI) and $\delta=-0.04^{+0.09}_{-0.10}$ (Keck). Additionally, EHT observations of Sgr A* estimated the angular diameter of the emission ring to be $\theta_{d}=(51.8\pm 2.3)\,\mu\text{as}$ [The Event Horizon Telescope Collaboration] et al. (2022a, b).
@@
For Sgr A*, we consider the averaged shadow angular diameter, derived from independent algorithms analyzing EHT observations. These studies suggest that $\theta_{d}$ for Sgr A* lies within the range $46.9\,\mu\text{as}$ to $50\,\mu\text{as}$. Including the $1\sigma$ uncertainty, this range extends from $41.7\,\mu\text{as}$ to $55.6\,\mu\text{as}$ [The Event Horizon Telescope Collaboration] et al. (2022b). In Fig. 8(b), we incorporate this confidence interval to compare the theoretical shadow diameter with the observational results, thereby constraining $A_{0}$. Our findings suggest that the LGT black hole aligns most closely with the EHT data when $A_{0}$ approaches the Schwarzschild limit, as $A_{0}$ remains nearly unity within its uncertainty range.

### v2

@@ V Shadow of the Black Hole
The photon sphere around a black hole is defined by unstable circular photon orbits that are significantly influenced by the gravitational field. Photons in these orbits either spiral into the black hole’s event horizon or escape to infinity, forming a luminous photon ring. This ring delineates the observable boundary of the black hole shadow for distant observers Synge (1966); Cunningham and Bardeen (1972); Bardeen et al. (1972); Luminet (1979). Historically, different terminologies such as escape cone Synge (1966), cone of gravitational radiation capture Zeldovich and Novikov (1971), optical appearance of black holes, and black hole image Bardeen (1973); Chandrasekhar (1998); Luminet (1979, 2018) have been utilized. However, the contemporary and widely adopted term black hole shadow, introduced by Falcke et al. Falcke et al. (2000), refers explicitly to the dark region surrounded by the photon ring Johannsen and Psaltis (2010); Johnson et al. (2020). In this section, we investigate the properties of the black hole shadow, focusing particularly on its dependence on the underlying model parameters, such as $A_{0}$, $m_{s}$, and $m_{p}$. Given the spherical symmetry of the spacetime considered, the analysis is simplified by restricting null geodesic trajectories to the equatorial plane ($\theta=\pi/2$). The metric in this case reduces to:
@@
$r_{\mathrm{sh}}=\frac{3\sqrt{3}A_{0}^{3}m_{s}}{8\pi m_{p}^{2}}.$ (81)
Another way to interpret the physical impact of the parameter $A_{0}$ and opens a direct path to confronting our model with observational data from the EHT [The Event Horizon Telescope Collaboration] et al. (2019a, 2021). We define the Schwarzschild mass as $M=m_{s}/(8\pi m_{p}^{2})$. The standard Schwarzschild photon sphere and shadow radii are $r_{\text{ph-Sch}}=3M$ and $r_{\text{sh-Sch}}=3\sqrt{3}M$, respectively. We can now re-express our results from Eqs. (79) and (81) in these terms. Photon sphere radius is obtained as
$r_{\mathrm{ph}}=\frac{3A_{0}^{2}m_{S}}{8\pi m_{P}^{2}}=3A_{0}^{2}M=A_{0}^{2}\,r_{\text{ph-Sch}},$ (82)
then we calculate black hole shadow radius
$r_{\mathrm{sh}}=\frac{3\sqrt{3}A_{0}^{3}m_{s}}{8\pi m_{p}^{2}}=3\sqrt{3}A_{0}^{3}M=A_{0}^{3}r_{\text{sh-Sch}}.$ (83)
The shadow size is directly modulated by $A_{0}^{3}$: if $A_{0}>1$, the shadow is larger than predicted by General Relativity, while for $A_{0}<1$, the reverse holds. Figures 3 highlight the sensitivity of the shadow radii to the parameter $A_{0}$. The shadow radius increases with the parameter $A_{0}$. This result underscores the potential observational impact of different model parameters in gravitational lensing phenomena.
This provides a powerful test for our model. The angular diameter of the shadow as seen from Earth is given by $\theta_{\text{sh}}=2r_{\text{sh}}/D$, where $D$ is the distance to the black hole. Using our derived expression, we can solve for $A_{0}$:
$\theta_{\text{sh}}=\frac{2A_{0}^{3}r_{\text{sh-Sch}}}{D}=\frac{6\sqrt{3}A_{0}^{3}M}{D}\quad\implies\quad A_{0}=\left(\frac{\theta_{\text{sh}}D}{6\sqrt{3}M}\right)^{1/3}$ (84)
We now apply this to the EHT observations for M87* and Sgr A*:
  1. 1.
    For M87*: Using the observational values of $\theta_{\text{sh}}=(42\pm 3)\,\mu\text{as}$, $M=6.5\times 10^{9}\,M_{\odot}$, and $D=16.8\,\text{Mpc}$, we find a constraint of $A_{0}\approx 1.04\pm 0.08$.
  2. 2.
    For Sagittarius A* (Sgr A*): Using $\theta_{\text{sh}}=(48.7\pm 7)\,\mu\text{as}$, $M=4.3\times 10^{6}\,M_{\odot}$, and $D\approx 8.2\,\text{kpc}$, we find a constraint of $A_{0}\approx 0.98\pm 0.14$.
Both results are remarkably consistent with General Relativity ($A_{0}=1$) within their respective uncertainties.
Figure 3: Shadow radius as functions of the parameter $A_{0}$, assuming $M=1$.
@@ VI Strong lensing by the black hole
$A(r_{0})\dot{t}_{0}^{2}=C(r_{0})\dot{\phi}_{0}^{2},$ (85)
Using the definition of the impact parameter in Eq. (75) at $r=r_{0}$, the function $R(r)$ in Eq. (76) can be rewritten as
$R(r)=\frac{A(r_{0})C(r)}{A(r)C(r_{0})}-1.$ (86)
We define the critical impact parameter as Bozza (2002)
$b_{c}\equiv b(r_{\mathrm{ph}})=\lim_{r_{0}\rightarrow r_{\mathrm{ph}}}\sqrt{\frac{C(r_{0})}{A(r_{0})}}.$ (87)
In Ref. Bozza (2002), the deflection angle of light rays passing at the radial distance $r_{0}$ from a static asymptotically flat black hole is given as
$\alpha(r_{0})=I(r_{0})-\pi=2\int_{r_{0}}^{\infty}\frac{\sqrt{B(r)}}{\sqrt{C(r)R(r)}}\,dr-\pi:=2\int_{r_{0}}^{\infty}f(r)\,dr-\pi,$ (88)
where $R(r)$ is given in Eq. (86). Introducing the variable $z\equiv 1-r_{0}/r$, the integral in Eq. (88) can be separated into the divergent part, $I_{D}(r_{0})$, and the regular part, $I_{R}(r_{0})$. The divergent part is expressed as Bozza (2002); Tsukamoto (2018b); Ali et al. (2025)
$I_{D}(r_{0})=\int_{0}^{1}f_{0}(z,r_{0})\,dz,$ (89)
@@
$x_{1}(r_{0})=\frac{1-A(r_{0})}{C(r_{0})A^{\prime}(r_{0})}\Big{[}C^{\prime}(r_{0})A(r_{0})-C(r_{0})A^{\prime}(r_{0})\Big{]},$ (91a)
$x_{2}(r_{0})=\frac{\bigl{(}1-A(r_{0})\bigr{)}^{2}}{2C(r_{0})^{2}A^{\prime}(r_{0})^{3}}\Big{[}2C(r_{0})C^{\prime}(r_{0})A^{\prime}(r_{0})^{2}+\bigl{(}C(r_{0})C^{\prime\prime}(r_{0})-2C^{\prime}(r_{0})^{2}\bigr{)}A(r_{0})A^{\prime}(r_{0})-C(r_{0})C^{\prime}(r_{0})A(r_{0})A^{\prime\prime}(r_{0})\Big{]}.$ (91b)
@@
$x_{1}(r_{0})=\frac{C(r_{\mathrm{ph}})r_{\mathrm{ph}}\mathfrak{L^{\prime}(r_{\mathrm{ph}})}}{B(r_{\mathrm{ph}})}\left(r_{0}-r_{\mathrm{ph}}\right)+\mathcal{O}(r_{0}-r_{\mathrm{ph}})^{2},$ (93)
@@
$b(r_{0})=b_{c}+\frac{1}{4}\sqrt{\frac{C(r_{\mathrm{ph}})}{A(r_{\mathrm{ph}})}}\,\mathfrak{L}^{\prime}(r_{\mathrm{ph}})\left(r_{0}-r_{\mathrm{ph}}\right)^{2}+\mathcal{O}(r_{0}-r_{\mathrm{ph}})^{2}.$ (94)
@@
$\lim_{r_{0}\rightarrow r_{\mathrm{ph}}}x_{1}(r_{0})=\lim_{b\rightarrow b_{c}}\frac{2C(r_{\mathrm{ph}})r_{\mathrm{ph}}\sqrt{\mathfrak{L}^{\prime}(r_{\mathrm{ph}})}}{B(r_{\mathrm{ph}})}\left(\frac{b}{b_{c}}-1\right)^{1/2}.$ (95)
@@
$I_{D}(b)=-\frac{r_{\mathrm{ph}}}{\sqrt{x_{2}(r_{\mathrm{ph}})}}\ln\left(\frac{b}{b_{c}}-1\right)+\frac{r_{\mathrm{ph}}}{\sqrt{x_{2}(r_{\mathrm{ph}})}}\ln\Bigl{(}r_{\mathrm{ph}}^{2}\mathfrak{L}^{\prime}(r_{\mathrm{ph}})\Bigr{)}+\mathcal{O}(b-b_{c}),$ (96)
@@
$x_{2}(r_{\mathrm{ph}})=\frac{C(r_{\mathrm{ph}})\bigl{(}1-A(r_{\mathrm{ph}})\bigr{)}^{2}\Bigl{[}C^{\prime\prime}(r_{\mathrm{ph}})A(r_{\mathrm{ph}})-C(r_{\mathrm{ph}})A^{\prime\prime}(r_{\mathrm{ph}})\Bigr{]}}{2A(r_{\mathrm{ph}})^{2}C^{\prime}(r_{\mathrm{ph}})}.$ (97)
@@
$\bar{a}=\sqrt{\frac{2B(r_{\mathrm{ph}})A(r_{\mathrm{ph}})}{C^{\prime\prime}(r_{\mathrm{ph}})A(r_{\mathrm{ph}})-C(r_{\mathrm{ph}})A^{\prime\prime}(r_{\mathrm{ph}})}},$ (100a)
$\bar{k}=\bar{a}\ln\Bigl{(}r_{\mathrm{ph}}^{2}\mathfrak{L}^{\prime}(r_{\mathrm{ph}})\Bigr{)}+I_{R}(r_{\mathrm{ph}})-\pi.$ (100b)
@@
$b_{c}=\sqrt{\frac{C(r_{\mathrm{ph}})}{A(r_{\mathrm{ph}})}}=\frac{3\sqrt{3}A_{0}^{3}m_{S}}{8\pi m_{P}^{2}},$ (101)
@@
Since this integral cannot be computed analytically for general $A_{0}$, we note that for the Schwarzschild black hole case ($A_{0}=\pm 1$), the integral evaluates to $I_{R}^{\mathrm{SBH}}\approx 0.9496$, which matches the result reported in Ref. Bozza (2002). More generally, as $A_{0}^{2}\to 1$, the values of $I_{R}$ approach this Schwarzschild black hole value (see Fig. 4).
@@
Additionally, for $A_{0}^{2}=1$, we find $\bar{k}=-0.40023$, which corresponds to the Schwarzschild black hole lensing result reported in Ref. Bozza (2002). Finally, by incorporating these relations, we present in Fig. 5 the $b$-profile of the strong deflection angle as a function of $A_{0}$.
@@ VI.1 The lens equation and the EHT constraints
Assuming that $D_{OL}\gg b_{c}$, the angular position of the $n$th relativistic image is given by Bozza (2002)
@@
To apply these theoretical results to astrophysical observations, we consider the supermassive black holes M87* and Sgr A*. The black hole M87* has a mass of $(6.5\pm 0.7)\times 10^{9}\,M_{\odot}$ and is located at a distance of $D_{OL}=16.8\,\mathrm{Mpc}$ from Earth [The Event Horizon Telescope Collaboration] et al. (2019a, b). Meanwhile, Sgr A* has a mass of $4^{+1.1}_{-0.6}\times 10^{6}\,M_{\odot}$ and is situated at a distance of $7.97\,\mathrm{kpc}$ from Earth [The Event Horizon Telescope Collaboration] et al. (2022). Substituting these parameters into Eq. (112), Fig. 6 illustrates the outermost Einstein rings ($n=1$) for M87* and Sgr A*, assuming they are black holes within the framework of the LGT. These rings are depicted in the celestial coordinate system of an observer on Earth, with coordinates $X$ and $Y$, and are computed for different values of $A_{0}$.
@@
A key quantity in strong lensing is the magnification of the $n$th relativistic image, defined as Virbhadra and Ellis (2000); Bozza (2002)
@@
Since the magnification is inversely proportional to $D_{OL}^{2}$, relativistic images are typically faint. The outermost image appears the brightest, while the brightness of higher-order images diminishes exponentially. However, in the limiting case where $\psi$ approaches zero, implying near-perfect alignment of the source and the lens, the images undergo significant magnification. It is noteworthy that while the outermost image, corresponding to $\theta_{1}$, remains distinct, higher-order images tend to cluster around $\theta_{\infty}\equiv\theta_{n}|_{n\to\infty}$ Bozza (2002), which is identified as $\theta_{\infty}=\theta_{c}$, representing the innermost relativistic image.
@@
while the second is the relative magnification between the outermost relativistic image and the collective group of inner relativistic images, expressed as Bozza (2002)
@@
In 2019, the EHT collaboration published the first horizon-scale image of the supermassive black hole M87*, providing strong observational evidence for black holes. Their analysis revealed that the compact emission region had an angular diameter of $\theta_{d}=(42\pm 3)\,\mu\text{as}$, along with a central flux depression exceeding a factor of $\gtrsim 10$, corresponding to the black hole’s shadow [The Event Horizon Telescope Collaboration] et al. (2019a, b). Subsequently, in 2022, the EHT collaboration captured an image of Sgr A*, the supermassive black hole at the center of the Milky Way. This image displayed a distinctive ring structure with an angular diameter of $\theta_{d}=(48.7\pm 7)\,\mu\text{as}$ and a deviation from the Schwarzschild shadow characterized by $\delta=-0.08^{+0.09}_{-0.09}$ (VLTI) and $\delta=-0.04^{+0.09}_{-0.10}$ (Keck). Additionally, EHT observations of Sgr A* estimated the angular diameter of the emission ring to be $\theta_{d}=(51.8\pm 2.3)\,\mu\text{as}$ [The Event Horizon Telescope Collaboration] et al. (2022).
@@
For Sgr A*, we consider the averaged shadow angular diameter, derived from independent algorithms analyzing EHT observations. These studies suggest that $\theta_{d}$ for Sgr A* lies within the range $46.9\,\mu\text{as}$ to $50\,\mu\text{as}$. Including the $1\sigma$ uncertainty, this range extends from $41.7\,\mu\text{as}$ to $55.6\,\mu\text{as}$ [The Event Horizon Telescope Collaboration] et al. (2022). In Fig. 7(b), we incorporate this confidence interval to compare the theoretical shadow diameter with the observational results, thereby constraining $A_{0}$. Our findings suggest that the LGT black hole aligns most closely with the EHT data when $A_{0}$ approaches the Schwarzschild limit, as $A_{0}$ remains nearly unity within its uncertainty range.